OpenStudy (anonymous):

The sum of 3 consecutive integers is 261. What is the middle number?

OpenStudy (alina123):

1st integer—x: x + x + 2 + x + 4 = 261 3x = 255 x = 85 2nd integer: = 85 + 2 = 87 3rd integer: = 85 + 4 = 89 Answer: 85, 87 & 89 are the 3 consecutive odd integers. Proof (sum): = 85 + 87 + 89 = 261

OpenStudy (anonymous):

thankyou